OpenStudy (anonymous):

what is the cube root of 64? how do you work it?

OpenStudy (anonymous):

64=2^6 or 64=4^3 Hope this helps

OpenStudy (anonymous):

How do did you get that? also it helps heaps thanks :)

OpenStudy (anonymous):

Well over 18 years of maths study helps :) The powers of 2 I know even if I just wake up from my afternoon nap

OpenStudy (anonymous):

factor out 64 64 = 8(8) = 2(4)(2)(4) = 4(4)(4) so, cube root of 64 = cube root of [4(4)(4)] = 4 hope this helps

OpenStudy (anonymous):

factor out 64 64 = 8(8) = 2(4)(2)(4) = 4(4)(4) so, cube root of 64 = cube root of [4(4)(4)] = 4 then 64=(4)^3 => cube root 64=4 hope this helps

OpenStudy (anonymous):

The proper way to do this question if you do not know that 64=2^6 is What are the divisors of 64. It is an even number so 2 is 64/2=32 32 is even 32/2=16 ..... 16/2=8 8/2=4 4/2=2 2/2=1 How many times did I do that 32,16,8,4,2,1 6 times So 64=2^6 6=2*3 64=2^(2*3)=4^3

OpenStudy (anonymous):

Thank yous so much. If you put down 2^6 would it matter, or do you have to find a number that is only multiplied 3 times..4x4x4 since its cube?

OpenStudy (anonymous):

The answer is 4: as the question asks for the cube root